OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). ?

OpenStudy (mertsj):

The slope of the given line is 2 so the slope of the perpendicular is -1/2 Since y = mx + b substitute all the known values and find b. y = 3, x = 6 and m = -1/2

OpenStudy (mertsj):